OpenStudy (anonymous):

Describe the positive integer values of n for which (x^2+\frac{1}{x})^n has a nonzero constant term. In terms of n, what is that constant term? (You can leave your answer in the form of a combination.)

OpenStudy (anonymous):

can you type the expression in equation editor or write it out on the draw please?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[\Big(x^2+\frac{1}{x}\Big)^n\]

OpenStudy (anonymous):

do a few... staring with n = 1 => no constant term you should start to see a pattern. binomial theorem says \[\left( a+b \right)^{n}= \left(\begin{matrix}n \\ 0\end{matrix}\right)a ^{n}b ^{0}+\left(\begin{matrix}n \\ 1\end{matrix}\right)a ^{n-1}b ^{1}+\left(\begin{matrix}n \\ 2\end{matrix}\right)a ^{n-2}b ^{2}+...+\left(\begin{matrix}n \\ r\end{matrix}\right)a ^{n-r}b ^{r}+...+\left(\begin{matrix}n \\ n\end{matrix}\right)a ^{n-n}b ^{n}\] in your case, you'll need this to have a constant term \[\left( x ^{2} \right)^{k}\left( \frac{ 1 }{x } \right)^{2k}=1\] so that the constant term will be \[\left(\begin{matrix}2k+k \\ 2k\end{matrix}\right)\] so n must be...

OpenStudy (anonymous):

does that make any sense? you gonna reply?

OpenStudy (anonymous):

uhh...

OpenStudy (anonymous):

uhh?

OpenStudy (anonymous):

still reading through it

OpenStudy (anonymous):

okay... if you have any questions or need any clarification... please let me know

OpenStudy (anonymous):

does that mean that it must be an even number?

OpenStudy (anonymous):

what must be an even number?

OpenStudy (anonymous):

the constant term

OpenStudy (anonymous):

no... 3 chose 2 isn't even

OpenStudy (anonymous):

2k+k = n so n must be what kind of number?

OpenStudy (anonymous):

btw, k = 1, 2, 3, ...

OpenStudy (anonymous):

an odd number

OpenStudy (anonymous):

nope... what's 2k + k?

OpenStudy (anonymous):

3k

OpenStudy (anonymous):

oh. so its a multiple of 3!!!

OpenStudy (anonymous):

I'm such an idiot

OpenStudy (anonymous):

it's okay

OpenStudy (anonymous):

why did I not see that? you're really smart pgpilot

OpenStudy (anonymous):

i wish, but thanks give yourself some cred as well and don't get down on you!

OpenStudy (anonymous):

ok thank you so much!!

OpenStudy (anonymous):

you are more than welcome!

OpenStudy (anonymous):

keep the good questions like that one coming... my brain needs the challenge!